What is uncertainty principle.
1
Expert's answer
2020-10-12T07:28:45-0400

Solution

Heisenberg Uncertainty principle:this principle states that when we measure position with absolute value there is found uncertainty in momentum, in other hand if we measure momentum with absolute value then there is found uncertainty in position.

The product of uncertainties in position "\Delta x" and momentum "\Delta p" is "\frac{h}{4\pi}" Or more.

This is represented as

"\Delta x. \Delta p \ge \frac{h}{4\pi}"